OpenStudy (anonymous):

A $7000 investment earns $1330 simple interest in two years. What is the annual interest rate? A. 1.90% B. 3.80% C. 9.50% D. 19.0%

jimthompson5910 (jim_thompson5910):

You'll use the formula I = P*r*t where the interest earned is I = 1330, the amount invested is P = 7000 and the time is t = 2 years. So, I = P*r*t 1330 = 7000*r*2 then you solve for r

jimthompson5910 (jim_thompson5910):

The value of r that you get will be in decimal form. Multiply that result by 100 to convert to percent form.

OpenStudy (anonymous):

so... i guess ill take the red pill.

OpenStudy (anonymous):

WELCOME TO THE MATRIX